5) A disk is initially spinning at 14 counter-clockwise. After 5 seconds, it is spinning clockwise and a piece of dust located 1 meter from the center of disk is moving at 7 m/s. How many revolutions does the disk make before it stopped spinning?
Answer #1

Apply rotational kinematic equation

wf- wi = alpha t

alpha = wf- wi/t = (7m/s/1 m) -(-14)/5 ( wf = v/r)

= 4.2 rad/s^2

del theta = wf^2 - wi^2/ 2*alpha

= 0- ( -14)^2/2 ( 4.2)

=23.33 rad ( rev/ 2pi)

=3.71 rev